1) Julia Willis will invest $24,900 today. She needs $63,826 in 24 years.
What annual interest rate must she earn? (Round answer to 0 decimal places, e.g. 7%.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

63,826=24900*(1+r/100)^24

(63,826/24900)^(1/24)=(1+r/100)

(1+r/100)=1.04

r=1.04-1

=4%